# Magic dynamics in many-body localized systems

### v1

@@ Abstract
Nonstabilizerness, also known as quantum magic, characterizes the beyond-Clifford operations needed to prepare a quantum state and constitutes an essential resource, alongside entanglement, for achieving quantum advantage. This work investigates how nonstabilizerness spreads under the dynamics of disordered quantum many-body systems. Using the $\ell$-bit model, a phenomenological model of many-body localization (MBL), we present an analytical description of the nonstabilizerness growth in MBL systems. We demonstrate that our analytical formulas describe the nonstabilizerness growth in strongly disordered quantum spin chains. Our findings establish a new facet of MBL phenomenology and identify the vital role of the disorder in slowing down the growth of the complexity of quantum states, important for our understanding of quantum advantage.
@@ Introduction.
Quantum state |$\ket{\psi}$ of $L$ qubits is specified by a state vector in $2^{L}$ dimensional Hilbert space [1]. The exponential growth of many-body Hilbert space implies that quantum states may become intractable for classical computers for sufficiently large $L$ [2], motivating the development of quantum simulators [3, 4, 5] and quantum computers [6, 7, 8, 9]. However, certain quantum states possess a structure that enables their efficient representation on classical computers. For instance, when |$\ket{\psi}$ is weakly entangled [10], it can be simulated at cost increasing polynomially with $L$ using tensor network approaches [11, 12, 13, 14, 15, 16]. Hence, extensive entanglement [17, 18] is necessary for quantum devices to reach computational advantage over classical computers [19]. Nevertheless, stabilizer states [20, 21], may host an extensive entanglement, and still be simulated with classical resources scaling polynomially with $L$ [22, 23]. Therefore, nonstabilizerness, commonly referred to as “magic”, quantifying the extent to which |$\ket{\psi}$ departs from the set of stabilizer states [24, 25, 26, 27, 28, 29], is a quantum resource [30] essential for characterizing the complexity of quantum states.
@@
Understanding the generation of magic resources in many-body systems is fundamental for assessing their classical simulability. Generically, quantum many-body systems prepared in an out-of-equilibrium state are expected to follow eigenstate thermalization hypothesis (ETH[31, 32, 33, 34] and to thermalize reaching an equilibrium state described by appropriate ensembles of statistical mechanics [35, 36, 37, 38]. Thermalization is accompanied by fast, ballistic [39, 40] or sub-ballistic [41, 42], growth of entanglement entropy, and a rapid saturation of nonstabilizerness measures [43, 44] to their maximal values [45]. The process of thermalization slows down in the presence of disorder [46, 47, 48, 49]. Sufficiently strong disorder leads to a phenomenon of the many-body localization (MBL[50, 51, 52, 53] which prevents the thermalization [54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68] at any experimentally relevant time scale [69, 70]. The absence of thermalization starkly affects the dynamics of MBL systems, leading to a logarithmic in time growth of entanglement entropy [71, 72, 73, 74, 75] and the memory of the initial state due to the presence of an emergent set of local integrals of motion, dubbed localized bits or $\ell$-bits [76, 56]. This raises the question: how does magic spread in MBL systems?
@@
Figure 1: Magic spreading in disordered quantum systems. In the absence of interactions (III), $\mathcal{M}_{2}$ is described solely by spin precession and saturates to $L\log_{2}(8/7)$. In the interacting case, slow spin dephasing leads to a power-law growth of $\mathcal{M}_{2}$. For the initial $X$-polarized state, magic grows as described by Eq. 5, c.f. (I), and saturates on a timescale much shorter than that required for the entanglement entropy to saturate. For a generic random product state, (II), the growth of nonstabilizerness is limited.
@@
In this work, we address this question by investigating the dynamics of nonstabilizerness in strongly disordered spin chains that exhibit MBL. Leveraging the $\ell$-bits, we develop a theoretical framework that describes how magic evolves under MBL dynamics. Specifically, we analyze the stabilizer Rényi entropy (SRE) [28] as a measure of magic, characterizing its growth in both non-interacting and interacting disorder localized systems. In non-interacting localized systems, the SRE remains limited, while in MBL systems, the slow spin dephasing leads to a power-law growth of nonstabilizerness before its eventual saturation. To corroborate our theoretical predictions, we perform numerical simulations on a microscopic MBL model, the disordered transverse-field Ising model (TFIM). We track the evolution of magic across different disorder strengths, initial states, and system sizes, showing remarkable agreement with the $\ell$-bits-based predictions. Additionally, we demonstrate that nonstabilizerness enables a clear distinction between MBL and ergodic regimes.
@@
where $L$ is the number of qubits, $k$ is the Rényi index, and $P$ is a Pauli string that belongs to the Pauli group $\mathcal{P}_{L}$. In particular, $\mathcal{M}_{1}$ is defined by the limit $k\to 1$ in (1), and $\mathcal{M}_{k}\geq 0$, with the equality holding if and only if $|\Psi\rangle$ is a stabilizer state [77, 78]. In our study, we fix $k=2$ and evaluate the SRE using the algorithm of [79], which allowed us to obtain numerically exact results for $L\leq 22$. One advantage of the SRE over many other proposed measures of magic [26] is that it allows an efficient computation even for a large $L$  [80, 81, 82, 83, 84, 85]. Moreover, $\mathcal{M}_{k}(|\Psi\rangle)$ is also experimentally measurable [85, 86, 87, 88, 89].
Quantifying nonstabilizerness and understanding how these resources grow is a current topic of interest in quantum many-body physics [27]. Recent works have addressed this question for ergodic many-body systems [45, 44], where initial state information is rapidly lost and $|\Psi\rangle$ behaves similarly to a random vector [34]. For example, in random unitary circuits, the SRE saturates to the Haar-random state value [90]
@@
^$\hat{\mathcal{H}}_{\mathrm{\ell-bit}}=\sum_{i}h_{i}\hat{\tau}_{i}^{z}+\sum_{i<j}J_{ij}\hat{\tau}_{i}^{z}\hat{\tau}_{j}^{z}+\sum_{i<j<k}J_{ijk}\hat{\tau}_{i}^{z}\hat{\tau}_{j}^{z}\hat{\tau}_{k}^{z}+...$ (3)
where $h_{i}$ are random on-site fields, $J_{ij...}$ are interaction terms that decay exponentially with the distance between the spins, and $\hat{\tau}^{z}$ are quasilocal operators that mutually commute [94]. The $\ell$-bit model, (3), captures many phenomenological properties of MBL, including eigenvalue statistics [95, 96], entanglement [97, 98, 99], and other aspects of the dynamics [100, 101, 102]. In the following, we utilize (3) to understand nonstabilizerness dynamics in an MBL system.
@@
[…]
for the initial $|\Psi_{X}^{+}\rangle$ state. In the limit of $t\rightarrow\infty$, this expression yields $\mathcal{M}_{2}\approx L\log_{2}(8/7)$, see Fig. 1. For a generic product initial state $|\Psi_{R}\rangle$, our numerical results show that the SRE rapidly saturates to a constant value not bigger than $\mathcal{M}_{2}(|\Psi\rangle)+L\log_{2}(8/7)$, see [94].
Figure 2: Dynamics of nonstabilizerness in disordered TFIM. Evolution of SRE for initial states (a) $|\Psi_{Z}\rangle$, (b) $|\Psi_{X}\rangle$, (c) $|\Psi_{Y}\rangle$, and (d) $|\Psi_{R}\rangle$ (see text). The results are for $L=16$ and averaged over $1000$ realizations, considering the product state close to the middle of the spectrum. To demonstrate the validity of (5), we performed a numerical fit at $W=5$ for all states. The saturation value $\mathcal{M}_{2}^{\mathrm{sat}}$ depends on the initial state: $\mathcal{M}_{2}^{\mathrm{sat}}\approx 7.3$ for $|\Psi_{Z}\rangle$, $\mathcal{M}_{2}^{\mathrm{sat}}\approx 11.53$ for $|\Psi_{R}\rangle$, while the SRE for $|\Psi_{X}\rangle$ and $|\Psi_{Y}\rangle$ states saturates to the Haar value $\mathcal{M}_{2}^{\mathrm{Haar}}$. The power-law growth exponents are $\beta\approx 0.16$ for $|\Psi_{Z}\rangle$, $\beta\approx 0.29$ for $|\Psi_{X}\rangle$, $\beta\approx 0.39$ for $|\Psi_{Y}\rangle$, and $\beta\approx 0.19$ for $|\Psi_{R}\rangle$. Similar behavior is obtained for other disorder strengths within the MBL regime.
@@
To understand the dynamics of SRE, it is essential to examine the impact of the spin dephasing terms on the time evolution of Pauli strings expectation values [103]. The expectation values of Pauli strings composed solely of $\hat{\tau}^{z}$ operators remain constant under the dynamics of ^$\hat{\mathcal{H}}_{\ell\mathrm{-bit}}$. On the other hand, Pauli strings containing $\hat{\tau}^{x}$ or $\hat{\tau}^{y}$ exhibit distinctly different behavior. The slow dephasing of the spins causes a power-law decay of the single-spin expectation values $|\hat{\tau}_{j}^{\alpha}|$, with $\alpha\in\{x,y\}$; for multi-spin observables, the situation depends on whether the spins corresponding to $\hat{\tau}^{\alpha}$ operators are entangled. If these spins are entangled, the expectation value decays in the same power-law fashion as for the single-spin observables. However, if these spins are not entangled, the expectation value decays as a product of individual single-spin observables and, therefore, decays much faster. Before the spins get entangled, the sum of all Pauli strings results in a stretched exponential behavior of the sum in (1), leading to a power-law growth of $\mathcal{M}_{2}$ with a certain exponent $\beta$. As the particles gradually become entangled, this exponent decreases, and the growth of the SRE slows down with time. This results in the following dynamics of SRE in the MBL regime
@@
where $\mathcal{M}_{2}^{\mathrm{sat}}$, $c$ and $\beta$ are constants dependent on the initial state. In particular, for the $X$-polarized initial state $|\Psi_{X}^{+}\rangle$, the saturation value $\mathcal{M}_{2}^{\mathrm{sat}}=\mathcal{M}_{2}^{\mathrm{Haar}}$ is the same as for the ergodic dynamics, $c=\ln(D)$, and $\beta=\xi\ln(2)$ is the localization length of $\ell$-bits [94], see the orange dashed line in Fig. 1. For a generic initial product state, the saturation value is smaller, $\mathcal{M}_{2}^{\mathrm{sat}}<\mathcal{M}_{2}^{\mathrm{Haar}}$ and $\beta<\xi\ln(2)$ [94].
@@
^$\hat{\mathcal{H}}_{\mathrm{TFIM}}=\sum_{i=1}^{L-1}J_{i,i+1}\hat{Z}_{i}\hat{Z}_{i+1}+\sum_{i=1}^{L}h_{i}\hat{Z}_{i}+g\sum_{i=1}^{L}\hat{X}_{i}$ (6)
@@
where $h_{i}\in[-W,W]$ are random on-site fields that are drawn from a uniform distribution, $g$ is the transverse field, and $J_{i,i+1}$ are the interactions between neighboring spins. Building on [104], we consider nearest-neighbor couplings drawn from a uniform distribution $J_{i,i+1}\in[0.8,1.2]$, and fix the transverse field at $g=1$. Mathematical arguments [104, 105] suggest that this model hosts an MBL phase for sufficiently large disorder strength $W$. Finite-size scaling analysis [66, 106] places the critical disorder threshold at $W_{c}\approx 3.5$. However, similar to the XXZ chain [65], the model exhibits finite-size drifts [107]. We study the quench dynamics of an initial state $|\Psi\rangle$ using the Chebyshev polynomial expansion [108, 69] up to $t=2\times 10^{4}J$. We consider chains of $L\in[8,20]$ spins, with results averaged over $1000$ disorder realizations. The chosen initial state $|\Psi\rangle$ is a product state with energy near the middle of the spectrum.
@@
In Fig. 2(a), we show the time evolution of a $Z$ polarized initial state $|\Psi_{Z}\rangle$. This state belongs to the computational basis and is a conventional choice in MBL studies. For weak disorder, $W\sim 1$, $\mathcal{M}_{2}$ quickly grows towards the Haar value in the weak disorder regime, consistent with the ergodic dynamics results [43, 44]. For strong disorder $W\sim 5$, $|\Psi_{Z}\rangle$ is close to an eigenstate of (6) and, therefore, SRE increases very slowly. In the MBL regime, the behavior of SRE is accurately captured by the phenomenological formula (5), with the saturation value $\mathcal{M}_{2}^{\mathrm{sat}}$ considerably smaller than $\mathcal{M}_{2}^{\mathrm{Haar}}$.
@@
In Fig. 2(b), we present the SRE evolution of initial $X$-polarized state $|\Psi_{X}\rangle$, a random state in the $X$ basis. The SRE, $\mathcal{M}_{2}$, initially grows rapidly before slowing down at longer times. Even for strong disorder, the SRE approaches the Haar value at late times. A similar trend is observed for a randomly selected state of $Y$ basis, $|\Psi_{Y}\rangle$, as shown in Fig. 2(c). This choice leads to the smallest contribution of the Pauli strings involving only $\hat{Z}_{i}$ operators, and, consequently, the saturation value in (5) is $\mathcal{M}_{2}^{\mathrm{Haar}}$.
Additionally, we examine product states constructed by random rotations on the Bloch sphere, denoted as $|\Psi_{R}\rangle$. Since such states cannot be constructed by Clifford gates, $\mathcal{M}_{2}\neq 0$ at $t=0$. However, this initial value is significantly below the Haar limit, as non-entangled states typically possess low nonstabilizerness [109]. Therefore, under Hamiltonian dynamics, the magic resources spread over time, consistently with (5), and the SRE growth is intermediate between the dynamics for the $|\Psi_{Z}\rangle$ and $|\Psi_{Y}\rangle$ states.
Collapsing of stabilizer entropy.
@@
We consider $|\Psi_{Y}\rangle$ as our initial state. As shown in Fig. 3 (a), both $\ell$-bit model and TFIM exhibit similar dynamics, with $\mathcal{M}_{2}(\mathcal{S})$ collapsing onto a single master curve without any fitting parameters. For an initial random product state $|\Psi_{R}\rangle$ state the $\ell$-bit model predictions are aligned with the TFIM results when a disorder-dependent rescaling function, $f(W)$, is introduced. In Fig. 3(b) we show $\mathcal{M}_{2}/f(W)$ for $|\Psi_{R}\rangle$, where $f(W)$ is found by minimizing deviations of TFIM results from the $\ell$-bit model predictions. Importantly, $f(W)$ is independent of system size $L$. These results demonstrate a close connection between the SRE and entanglement growth in the MBL regime.
Finally, we investigate whether $\mathcal{M}_{2}$ can serve as a diagnostic for the crossover between ETH and MBL regimes. The SRE has been linked to phase transitions in many ground-state problems [81, 110, 111, 112, 113] and, therefore, one may expect that it is sensitive to the ETH-MBL crossover observed for $L\approx 20$ at $W_{c}\sim 3.5$ [66]. We consider time evolution of $|\Psi_{Z}\rangle$ and compute $\Delta\mathcal{M}_{2}=\mathcal{M}_{2}^{\mathrm{Haar}}-\mathcal{M}_{2}$ for different size systems $L$ at the longest available time ($t=2\times 10^{4}$), as shown in Fig. (4)(a). For weak disorder strength, $\Delta\mathcal{M}_{2}\rightarrow 0$ as $L$ increases, as expected for the ETH regime. On the other hand, $\Delta\mathcal{M}_{2}$ grows linearly with $L$ for strong disorder. The crossover between these two regimes occurs near $W_{c}\approx 3.5$, showing how the SRE distinguishes the ergodic and non-ergodic dynamical regimes.
Figure 4: Nonstabilizerness across the ETH-MBL crossover in TFIM. (a) Deviation from the Haar value $\Delta\mathcal{M}_{2}$ as a function of system size $L$ and for different disorder strengths. (b) $\Delta\mathcal{M}_{2}$ as a function of disorder strength $W$ at the latest time available for the $|\Psi_{Z}\rangle$ distinguishes the ETH and MBL regimes.
@@
In this work, we have explored the dynamics of the nonstabilizerness in disordered many-body systems exhibiting MBL. By developing a theoretical framework founded on an $\ell$-bit phenomenology, we have shown that magic spreading in MBL systems is fundamentally constrained by the slow dynamics characteristic of the MBL regime. Unlike ergodic systems, where SRE rapidly saturates to its maximal value, MBL systems display a much slower power-law relaxation towards the saturation value, which exhibits a strong variability with respect to the choice of initial state. Through numerical simulations of the disordered TFIM, we have verified our analytical formulas for the SRE growth and demonstrated a strong connection between the SRE and entanglement entropy grwoth in the MBL regime. In particular, our findings show that the disorder suppresses the rapid spread of magic resources, and interactions play a crucial role in enabling its slow but sustained growth. Furthermore, we have shown that SRE can serve as a diagnostic tool distinguishing ergodic and non-ergodic dynamics.

### v2

@@ Abstract
Nonstabilizerness, also known as quantum magic, quantifies the deviation of quantum states from stabilizer states, capturing the complexity necessary for quantum computational advantage. In this study, we investigate the dynamics of quantum magic in disordered many-body localized (MBL) systems using the stabilizer Rényi entropy (SRE). Leveraging a phenomenological description based on the $\ell$-bit model, we analytically and numerically demonstrate that interactions profoundly influence magic spreading, inducing a power-law growth of SRE that markedly contrasts with the rapid saturation observed in ergodic systems. We validate our theoretical predictions through numerical simulations of the disordered transverse-field Ising model, showing excellent agreement across various disorder strengths, system sizes, and initial states. Additionally, we uncover a universal relationship between SRE and entanglement entropy, revealing their common scaling in the MBL regime independent of disorder strength and system size. Our results offer critical insights into the interplay of disorder, interactions, and complexity in quantum many-body systems.
@@ Introduction.
[…]
@@
Understanding the generation of magic resources in many-body systems is fundamental for assessing their classical simulability. Generically, quantum many-body systems prepared in an out-of-equilibrium state are expected to follow the eigenstate thermalization hypothesis (ETH[32, 33, 34, 35] and to thermalize reaching an equilibrium state described by appropriate ensembles of statistical mechanics [36, 37, 38, 39]. Thermalization is accompanied by fast, ballistic [40, 41] or sub-ballistic [42, 43], growth of entanglement entropy, and a rapid saturation of nonstabilizerness measures [44, 45] to their maximal values [46]. The process of thermalization slows down in the presence of disorder [47, 48, 49, 50]. Sufficiently strong disorder leads to a phenomenon of many-body localization (MBL[51, 52, 53, 54] which prevents the thermalization [55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69] at any experimentally relevant time scale [70, 71]. The absence of thermalization starkly affects the dynamics of MBL systems, leading to a logarithmic in time growth of entanglement entropy [72, 73, 74, 75, 76] and the memory of the initial state due to the presence of an emergent set of local integrals of motion, dubbed localized bits or $\ell$-bits [77, 57]. This raises the question: how does nonstabilizerness spread in MBL systems?
@@
Figure 1: Magic spreading in disordered quantum systems. In the non-interacting case (III, red dashed line), $\mathcal{M}_{2}$ is analytically tractable (see [78]). It saturates rapidly to a finite value due to the absence of spin dephasing. In the MBL regime, dephasing induces a power-law growth of $\mathcal{M}_{2}$ toward a saturation value described by Eq. (5). For the initial X-polarized state (I, orange dashed line), magic grows rapidly and saturates to the Haar value. For a generic random product state (II, orange dashed line), the growth is slower and saturates at a lower value, revealing the dependence of magic spreading on the choice of initial state.
@@
In this work, we address this question by investigating the dynamics of nonstabilizerness in strongly disordered spin chains that exhibit MBL. Leveraging the $\ell$-bits, we develop a theoretical framework that describes how nonstabilizerness evolves under MBL dynamics. Specifically, we analyze the stabilizer Rényi entropy (SRE) [28] as a measure of magic, characterizing its growth in both non-interacting and interacting disorder localized systems. In non-interacting localized systems, the SRE remains limited, while in MBL systems, the slow spin dephasing leads to a power-law growth of nonstabilizerness before its eventual saturation. To corroborate our theoretical predictions, we perform numerical simulations on a microscopic MBL model, the disordered transverse-field Ising model (TFIM). We track the evolution of magic across different disorder strengths, initial states, and system sizes, showing remarkable agreement with the $\ell$-bits-based predictions. Additionally, we uncover a universal relationship between the stabilizer Rényi entropy and entanglement entropy, highlighting that their dynamics collapse onto a single curve in the MBL regime, independent of disorder strength and system size.
@@
where $L$ is the number of qubits, $k$ is the Rényi index, and $P$ is a Pauli string that belongs to the Pauli group $\mathcal{P}_{L}$. In particular, $\mathcal{M}_{1}$ is defined by the limit $k\to 1$ in (1), and $\mathcal{M}_{k}\geq 0$, with the equality holding if and only if $|\Psi\rangle$ is a stabilizer state [79, 80]. In our study, we fix $k=2$ and evaluate the SRE using the algorithm of [81], which allowed us to obtain numerically exact results for $L\leq 22$. One advantage of the SRE over many other proposed measures of magic [26] is that it allows an efficient computation even for a large $L$  [82, 83, 84, 85, 86, 87, 88]. Moreover, $\mathcal{M}_{k}(|\Psi\rangle)$ is also experimentally measurable [87, 89, 90, 91, 92].
Quantifying nonstabilizerness and understanding how these resources grow is a current topic of interest in quantum many-body physics [27]. Recent works have addressed this question for ergodic many-body systems [44, 46, 93, 94, 45, 95], where initial state information is rapidly lost and $|\Psi\rangle$ behaves similarly to a random vector [35]. For example, in random unitary circuits, the SRE saturates to the Haar-random state value [46]
@@
$\hat{\mathcal{H}}_{\mathrm{\ell-bit}}=\sum_{i}h_{i}\hat{\tau}_{i}^{z}+\sum_{i<j}J_{ij}\hat{\tau}_{i}^{z}\hat{\tau}_{j}^{z}+\sum_{i<j<k}J_{ijk}\hat{\tau}_{i}^{z}\hat{\tau}_{j}^{z}\hat{\tau}_{k}^{z}+$... (3)
where $h_{i}$ are random on-site fields drawn uniformly from [W,W][-W,W], $J_{ij...}$ are interaction terms that decay exponentially with the distance between the spins, and $\hat{\tau}^{z}$ are quasilocal operators that mutually commute [78]. The $\ell$-bit model, (3), captures many phenomenological properties of MBL, including eigenvalue statistics [98, 99], entanglement [100, 101, 102], and other aspects of the dynamics [103, 104, 105]. In the following, we utilize (3) to understand nonstabilizerness dynamics in an MBL system.
@@
$\mathcal{M}_{2}=-\frac{L}{W}\int_{0}^{W}dh\log_{2}\bigg{[}1-\frac{1}{4}\sin^{2}(4ht)\bigg{]}$ (4)
for the initial $|\Psi_{X}^{+}\rangle$ state. In the limit of $t\rightarrow\infty$, this expression yields $\mathcal{M}_{2}\approx L\log_{2}(8/7)$, shown by (III) red dashed line in Fig. 1. For a generic product initial state $|\Psi_{R}\rangle$, our numerical results show that the SRE rapidly saturates to a constant value not bigger than $\mathcal{M}_{2}(|\Psi_{R}\rangle)+L\log_{2}(8/7)$ (see [78] for an analytical derivation for arbitrary initial states)
Figure 2: Dynamics of nonstabilizerness in disordered TFIM. Evolution of SRE for initial states (a) $|\Psi_{Z}\rangle$, (b) $|\Psi_{X}\rangle$, (c) $|\Psi_{Y}\rangle$, and (d) $|\Psi_{R}\rangle$ (see text). The results are for $L=16$ and averaged over 10001000 realizations, considering the product state close to the middle of the spectrum. To demonstrate the validity of (5), we performed a numerical fit at $W=5$ for all states. The saturation value $\mathcal{M}_{2}^{\mathrm{sat}}$ depends on the initial state: $\mathcal{M}_{2}^{\mathrm{sat}}\approx 7.3$ for $|\Psi_{Z}\rangle$, $\mathcal{M}_{2}^{\mathrm{sat}}\approx 11.53$ for $|\Psi_{R}\rangle$, while the SRE for $|\Psi_{X}\rangle$ and $|\Psi_{Y}\rangle$ states saturates to the Haar value $\mathcal{M}_{2}^{\mathrm{Haar}}$. The power-law growth exponents are $\beta\approx 0.16$ for $|\Psi_{Z}\rangle$, $\beta\approx 0.29$ for $|\Psi_{X}\rangle$, $\beta\approx 0.39$ for $|\Psi_{Y}\rangle$, and $\beta\approx 0.19$ for $|\Psi_{R}\rangle$. Similar behavior is obtained for other disorder strengths within the MBL regime.
@@
To understand the dynamics of SRE, it is essential to examine the impact of the spin dephasing terms on the time evolution of Pauli strings expectation values [100]. The expectation values of Pauli strings composed solely of $\hat{\tau}^{z}$ operators remain constant under the dynamics of $\hat{\mathcal{H}}_{\ell\mathrm{-bit}}$. On the other hand, Pauli strings containing $\hat{\tau}^{x}$ or $\hat{\tau}^{y}$ exhibit distinctly different behavior. The slow dephasing of the spins causes a power-law decay of the single-spin expectation values $|\hat{\tau}_{j}^{\alpha}|$, with $\alpha\in\{x,y\}$; for multi-spin observables, the situation depends on whether the spins corresponding to $\hat{\tau}^{\alpha}$ operators are entangled. If these spins are entangled, the expectation value decays in the same power-law fashion as for the single-spin observables. However, if these spins are not entangled, the expectation value decays as a product of individual single-spin observables and, therefore, decays much faster. Before the spins get entangled, the sum of all Pauli strings results in a stretched exponential behavior of the sum in Eq. (1), leading to a power-law growth of $\mathcal{M}_{2}$ with a certain exponent $\beta$. As the particles gradually become entangled, this exponent decreases, and the growth of the SRE slows down with time. This results in the following dynamics of SRE in the MBL regime
@@
where $\mathcal{M}_{2}^{\mathrm{sat}}$, $c$ and $\beta$ are constants dependent on the initial state. In particular, for the $X$-polarized initial state $|\Psi_{X}^{+}\rangle$, the saturation value $\mathcal{M}_{2}^{\mathrm{sat}}=\mathcal{M}_{2}^{\mathrm{Haar}}$ is the same as for the ergodic dynamics, and $\beta=\beta^{\prime}\ln(2)$ is the fastest exponent for all possible initial configurations. Its dynamics is illustrated by the (I) orange dashed line in Fig. 1 (further discussion can be found in the End Matter section). Moreover, for a generic initial product state, the saturation value is smaller, $\mathcal{M}_{2}^{\mathrm{sat}}<\mathcal{M}_{2}^{\mathrm{Haar}}$ and $\beta<\beta^{\prime}\ln(2)$, as depicted in (II) orange dashed line in Fig. 1.
@@
$\hat{\mathcal{H}}_{\mathrm{TFIM}}=\sum_{i=1}^{L-1}J_{i,i+1}\hat{Z}_{i}\hat{Z}_{i+1}+\sum_{i=1}^{L}h_{i}\hat{Z}_{i}+g\sum_{i=1}^{L}\hat{X}_{i}$ (6)
@@
where $h_{i}\in[-W,W]$ are random on-site fields that are drawn from a uniform distribution, $g$ is the transverse field, and $J_{i,i+1}$ are the interactions between neighboring spins. Building on [106], we consider nearest-neighbor couplings drawn from a uniform distribution $J_{i,i+1}\in[0.8,1.2]$, and fix the transverse field at $g=1$. Mathematical arguments [106, 107] suggest that this model hosts an MBL phase for sufficiently large disorder strength $W$. Finite-size scaling analysis [67, 108] places the critical disorder threshold at $W_{c}\approx 3.5$. However, similar to the XXZ chain [66], the model exhibits finite-size drifts [109]. We study the quench dynamics of an initial state $|\Psi\rangle$ using the Chebyshev polynomial expansion [110, 70] up to $t=2\times 10^{4}J$. We consider chains of $L\in[8,20]$ spins, with results averaged over $1000$ disorder realizations. The chosen initial state $|\Psi\gamma\rangle$ is a product state in the $\gamma$ basis ($\gamma\in\{X,Y,Z\}$), with each qubit randomly assigned as $|\pm\rangle_{\gamma}$, and the total energy of $|\Psi\gamma\rangle$ is close to the middle of the spectrum.
@@
In Fig. 2(a), we show the time evolution of a $Z$ polarized initial state $|\Psi_{Z}\rangle$. For weak disorder, $W\sim 1$, $\mathcal{M}_{2}$ quickly grows towards the Haar value in the weak disorder regime, consistent with the ergodic dynamics results [44, 45]. For strong disorder $W\gtrsim 5$, $|\Psi_{Z}\rangle$ is close to an eigenstate of (6) and, therefore, SRE increases very slowly. In the MBL regime, the behavior of SRE is accurately captured by the phenomenological formula (5) (red dashed line in the Fig. 2(a)), with the saturation value $\mathcal{M}_{2}^{\mathrm{sat}}$ considerably smaller than $\mathcal{M}_{2}^{\mathrm{Haar}}$.
In Fig. 2(b), we present the SRE evolution of the initial $X$-polarized state $|\Psi_{X}\rangle$. The SRE, $\mathcal{M}_{2}$, initially grows rapidly before slowing down at longer times. Even for strong disorder, the SRE approaches the Haar value at late times. A similar trend is observed for $|\Psi_{Y}\rangle$, as shown in Fig. 2(c). Additionally, we examine product states constructed by random rotations on the Bloch sphere, denoted as $|\Psi_{R}\rangle$. Since such states cannot be constructed by Clifford gates, $\mathcal{M}_{2}\neq 0$ at $t=0$. However, this initial value is significantly below the Haar limit, as the SRE of product states is limited to $\mathcal{M}_{2}=L\log_{2}(4/3)$ [28]. Therefore, under Hamiltonian dynamics, the magic resources spread over time, consistent with (5), and the SRE growth is intermediate between the dynamics for the $|\Psi_{Z}\rangle$ and $|\Psi_{Y}\rangle$ states.
Dependence of the initial state.
To explain the dependence of $\mathcal{M}_{2}^{\mathrm{sat}}$ on the choice of the initial state, we revisit the $\ell$-bit model. Starting from a random product state in the computational basis $|i^{\prime}\rangle$, we prepare the state
Figure 3: Initial state dependence of SRE and weight of Z gates in MBL regime. (a) SRE dynamics in the $\ell$-bit model for different choices of $\alpha$. (b) Power-law exponent $\beta$ characterizing the growth of $\mathcal{M}_{2}$ as a function of $\alpha$; time evolution of $\mathcal{W}_{Z}$ for different initial states for the (c) $\ell$-bit model and (d) TFIM. The saturation value of $\mathcal{M}_{2}$ depends on the degree of localization of the initial state in the $\ell$-bit basis and it is intrinsically connected to $\mathcal{W}_{Z}$.
$|\Psi_{H}\rangle=\frac{1}{\sqrt{Z}}\sum_{i=1}^{D}e^{-\alpha{d}(i,i^{\prime})}|i\rangle,\quad Z=\sum_{i=1}^{D}e^{-2\alpha d(i,i^{\prime})}$ (7)
where $d(i,i^{\prime})$ is the Hamming distance between two states and $\alpha$ is a parameter that controls the degree of localization in the computational basis. In Fig. 3(a), we show that the dynamics of $\mathcal{M}_{2}$ slows down as $\alpha$ increases, i.e. when the initial state becomes more localized in the computational basis. Saturation values $\mathcal{M}_{2}^{\mathrm{sat}}$ also critically depend on the value of $\alpha$. The power-law exponent $\beta$, as illustrated in Fig. 3(b), decreases monotonically as $\alpha$ increases, practically vanishing for $\alpha\geq 2$.
The reason for different values of $\mathcal{M}_{2}^{\mathrm{sat}}$ becomes clear if we write $\mathcal{M}_{2}=-\log_{2}(\mathcal{W}_{Z}+c)$, where $\mathcal{W}_{Z}$ is the contribution from the Pauli subgroup $\mathcal{P}_{IZ}$ containing only $\hat{I}$ and $\hat{Z}$ gates, whereas $c$ takes into account the remaining strings. It is easy to show (see the End Matter) that $\mathcal{W}_{Z}$ can be written as
$\mathcal{W}_{Z}(t)=\sum_{u,v,k\in\{0,1\}^{L}}|c_{u}(t)|^{2}|c_{v}(t)|^{2}|c_{k}(t)|^{2}|c_{u\oplus v\oplus k}(t)|^{2},$ (8)
As mentioned before, Pauli strings from $\mathcal{P}_{IZ}$ remain frozen during the dynamics of (3), so $\mathcal{W}_{Z}$ is a constant fixed by the initial state, as shown in Fig. 3(c). Similar behavior is obtained for the TFIM deep in the MBL regime ($W=8$), where $\tau_{i}^{z}\approx Z_{i}$. As illustrated in Fig. 3(d), the choice of $|\Psi_{Y}\rangle$ in the TFIM leads to $W_{Z}\approx 1/D$, in agreement with the completely delocalized case in the $\ell$-bit. On the other hand, the choice of $|\Psi_{Z}\rangle$ leads to a very slow dynamics of $\mathcal{W}_{Z}$, showing that this state is close to an eigenstate of (3) (see the End Matter for further discussion).
Collapsing of stabilizer entropy.
@@
We consider $|\Psi_{Y}\rangle$ as our initial state. As shown in Fig. 4 (a), both $\ell$-bit model and TFIM exhibit similar dynamics, with $\mathcal{M}_{2}(\mathcal{S})$ collapsing onto a single master curve without any scaling parameters. For an initial random product state $|\Psi_{R}\rangle$ state the $\ell$-bit model predictions are aligned with the TFIM results when a disorder-dependent rescaling function, $f(W)$, is introduced. In Fig. 4(b) we show $\mathcal{M}_{2}/f(W)$ for $|\Psi_{R}\rangle$, where $f(W)$ is found by minimizing deviations of TFIM results from the $\ell$-bit model predictions. Importantly, $f(W)$ is independent of system size $L$. These results demonstrate a close connection between the SRE and the growth of entanglement in the MBL regime.
@@
In this work, we have explored the dynamics of the nonstabilizerness in disordered many-body systems exhibiting MBL. By developing a theoretical framework founded on an $\ell$-bit phenomenology, we have shown that magic spreading in MBL systems is fundamentally constrained by the slow dynamics characteristic of the MBL regime. Unlike ergodic systems, where SRE rapidly saturates to its maximal value, MBL systems display a much slower power-law relaxation towards the saturation value, which exhibits a strong variability with respect to the choice of initial state. Through numerical simulations of the disordered TFIM, we have verified our phenomenological formulas for the SRE growth and demonstrated a strong connection between the SRE and entanglement entropy growth in the MBL regime. In particular, our findings show that the disorder suppresses the rapid spread of magic resources, and interactions play a crucial role in enabling its slow but sustained growth.